OpenStudy (anonymous):

A rectangular box is twice as long as it is wide. The height of the box is 3 feet less than the width. If the box is x feet wide, what polynomial represents its volume in cubic feet? A. 4x-3 B. 2x^3+6x^2 C. 2x^2-6x D. 2x^3-6x^2

OpenStudy (lovelyharmonics):

okay so x=width of box. length of box = 2 * x. Height of box = x - 3. Volume of box = length * width * height = 2x * x * (x-3) = 2x^2 * (x-3) = 2x^3 - 6x^2

OpenStudy (anonymous):

Hmmm

OpenStudy (anonymous):

I know its not A or C. So it between B and D. So, by just solving it together. I wanna say D

OpenStudy (lovelyharmonics):

its d c: i already put that

OpenStudy (anonymous):

Oh. Well Ima blonde xD

OpenStudy (lovelyharmonics):

i think my bodies secretly blond, my hair just dosent know it yet c:

OpenStudy (anonymous):

Lmao thats what I say xD This is the next question at least the file

OpenStudy (lovelyharmonics):

can you like attach it in a picture .-.cuz my computers being stupid

OpenStudy (anonymous):

I don't know how DX Most I know is email

OpenStudy (lovelyharmonics):

wait i think i got it c:

OpenStudy (anonymous):

Okay

OpenStudy (lovelyharmonics):

c cuz the ticket price would be 25 since the maximum income would be 625

OpenStudy (anonymous):

Are you kidding me! -tosses Algebra into the sun- There it no longer exist

OpenStudy (lovelyharmonics):

xD im taking online pre calc and i just google the questions, cuz 80% of the time someone else has already asked the same question on here or some other website like i just found that answer on yahoo XD